OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = 1,024 and a4 = –16?

OpenStudy (anonymous):

Well, you can use the formula: An = A1*r^(n-1) where An is the nth term of the sequence, A1 is the first term of the sequence, and r is the common ratio. First you have to figure out what the common ratio is. A4 = 1024*r^(3) = -16 solve for r

OpenStudy (anonymous):

Once you know r, plug that into the original formula and solve for A7

OpenStudy (anonymous):

Give it a try and ask me any questions that come up.